OpenStudy (anonymous):

Medals Given and fans Describe a situation that can be expressed with a two-step equation. Write the equation, solve it, and check the solution, showing all your work.

OpenStudy (anonymous):

@justnick09

OpenStudy (justnick09):

idk

OpenStudy (anonymous):

The late fee on overdue books is $1.50 per day. I already owe $12. How many days will it be until I owe $21 ? Equation to solve 12 + 1.5t = 21 -12 -12 1.5 t = 21 divide both sides by 1.5 t = 14 days